OpenStudy (anonymous):

The lengths of the eggs of a species of bird are roughly normally distributed, with a mean of 32 mm and an SD of 1.2 mm.Approximately what is the 99th percentile of the lengths? Is it posible to have some help?Thank you

Directrix (directrix):

To get the z-score for the 99th percentile, I use an online normal distribution table. Reading in the table from its interior out, I got z to be 2.33. 2.33 = (x - 32)/1.2 where x is the egg length at the 99th percentile position in the data. Cross multiply and solve for x (x - 32) = 2.33 * 1.2

Directrix (directrix):

I'll add a section about how to read the normal distribution area table for those like me who don't have a calculator handy.

OpenStudy (amistre64):

hmmm, same exact question was asked by another user ... and of course i gave a very similar answer :)

Directrix (directrix):

That question had the same bird data but a difference task. I think it was an interval containing 50% of the data problem. @amistre64

Directrix (directrix):

How to read a z-score from a table of areas under the standard normal distribution curve.

OpenStudy (amistre64):

thats the one :)